OpenStudy (anonymous):

Give an equation of the line through the points (3, –1) and (0, 6).

OpenStudy (anonymous):

I came up with y=-7/3x+8

OpenStudy (jdoe0001):

hmm what did you get for the slope?

OpenStudy (anonymous):

-7/3

OpenStudy (jdoe0001):

hmm ok so slope is 7/3 right and let's say let us use the 2nd point, (0, 6) \(\bf y-6=\frac{7}{3}(x-0)\implies y-6=\frac{7}{3}x\implies y=\frac{7}{3}x+6\)

OpenStudy (anonymous):

OK this is what i did y-(-1)=-7/3(x-3) y-(-1)=-7/3x+7 y=-7/3x+8 What did I do wrong?

OpenStudy (jdoe0001):

actually... should be negative 7/3 shoot lemme redo that

OpenStudy (anonymous):

Alright (-:

OpenStudy (jdoe0001):

anyhow \(\bf y-6=-\frac{7}{3}(x-0)\implies y-6=-\frac{7}{3}x\implies y=-\frac{7}{3}x+6\) in your case using the 1st point y-(-1)=-7/3(x-3) y-(-1)=-7/3x+7 -----> y+1=-7/3+7 => y= -7/3 +7 -1 y=-7/3x+8

OpenStudy (anonymous):

Alright so I solved it correct?

OpenStudy (jdoe0001):

well lemme do it using the 1st point \(\bf y-({\color{brown}{ -1}})=-\frac{7}{3}(x-3)\implies y{\color{brown}{ +1}}=-\frac{7}{3}x+7\implies y=-\frac{7}{3}x+7{\color{brown}{ -1}}\)

OpenStudy (anonymous):

OHHH OK it's y=-7/3x+6! Thank you!!!

OpenStudy (jdoe0001):

yw